Can you please help,Von is thrice as old as his brother Jon. Four years ago, he was 4 years less than 5 times Jon’s age at that time. How old are they now?

Answer by Theo(13342)

v = von's age
j = jon's age.
v = 3j is your first equation.
v-4 = 5(j-4)-4 is your second equation.
you have 2 equations that need to be solved simultaneously.
they are:
v = 3j
v-4 = 5(j-4)-4
simplify the second equation to get:
v-4 = 5j-20-4 which simplifies further to:
v-4 = 5j-24
add 4 to both sides of this equation to get:
v = 5j-20
your 2 equations are now:
v = 3j
v = 5j-20
since both expressions on the right side of each equation are equal to v, you can set these expressions equal to each other to get:
3j = 5j-20
subtract 5j from both sides of this equation to get:
-2s = -20
divide both sides of this equation by -2 to get:
j = 10
your first equation was:
v = 3j
this makes v = 30
your second equation was:
v-4 = 5(j-4)-4
replace v with 30 and j with 10 to get:
30-4 = 5(10-4)-4 which simplifies to:
26 = 50-20-4 which simplifies further to:
26 = 30-4 which simplifies further to:
26 = 26
the values of 30 for v and 10 for j are good.